6x - 5 = 20
I keep getting stuck on this when I try to solve it. Please show your work so I can understand how to solve it.
Thank You

Answer by checkley77(12844)

6x-5=20
6x=20+5
6x=25
x=25/6 or 4 1/6 or 4.16667 ans.